A mass of 620 g stretches a spring 4 cm. If the spring-mass system is placed in a medium that imparts a viscous force of 4 N when the speed of the mass is 9 cm/sec, find ?, the damping coefficient for this system. Use ?=9.8 m/sec2.
Answer #1

Here Viscous force = cv = 4

c = 4/v = 4/9e-2 = 44.44

  Damping coefficient  ? = c/2m = 44.44/(2*0.620) = 35.84 s^-1 Answer